Looking for the value of c when the system of equations is inconsistent (parallel lines).
2x+cy=11
5x-7y=5
Thanks!

Answer by rothauserc(4718)

parallel lines have the same slope, so let's put each equation in standard form, y = mx +b where m is the slope and b is the y intercept
2x + cy = 11
cy = -2x + 11
y = -2x/c + 11/c
5x - 7y = 5
-7y = -5x + 5
y = 5x/7 - 5/7
now set the two slopes equal to each other
5/7 = -2/c
5c = -14
c = -14/5